Solve each proportion
1.13/49=26/7x
2. b+1/b-1=5/6
3. b-1/b-1=5/6
1
Expert's answer
2020-10-20T12:26:44-0400

"1. \frac{13}{49} = \frac{26}{7x}\\
\textsf{Dividing both sides by} \frac{13}{7},\\

\frac{1}{7} = \frac{2}{x}\\
x = 14\\
2. \frac{b + 1}{b - 1} = \frac{5}{6}\\
\textsf{Multiplying both sides by 6(b-1)},\\
6b + 6 = 5b -5\\
6b - 5b = -5-6\\
b = -11\\
3.\frac{b-1}{b+1} = \frac{5}{6}\\
\textsf{Multiplying both sides by 6(b+1)}\\
6b-6 = 5b+5\\
6b-5b = 5+6\\
b=11"